Fermi National Laboratory
 

Operation Summary for June 16, 2010 - June 18, 2010

 

Wednesday June 16

The day shift began with TeV personnel conducting orbit corrections, with Operations monitoring the Recycler stash (215.0E10), with the Antiproton Source (Pbar) stacking (7.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

During the day shift, Operations established HEP store 7897 at 10:37 AM with an average initial luminosity (AIL) of 258E30.

During the evening shift, Operations monitored the accelerator complex.

Thursday June 17

The midnight shift began with Operations monitoring store 7897 (55.22E30) and the stash (346.6E10), with Pbar stacking (22.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7897 at 2:30 AM, put the TeV into shot setup, and established store 7898 at 4:15 AM with an AIL of 352E30.

During the day shift, Operations disabled the NuMI beam at 10:30 AM to allow NuMI personnel to investigate an inconsistent output from the horn power supply. Operations resumed sending beam to NuMI at 12:49 PM, but the horn tripped off at 1:37 PM and would not reset; EE Support personnel found a leaking LCW hose in the horn power supply. EE Support personnel successful hipotted HP3DS at 1 PM, but the power supply would not turn on; their investigation discovered all three 13.8kV fuses had blown. HP3DS passed a critical test at 4 PM, but again would not turn on. EE Support continued their investigations.

During the evening shift, Meson FTTBT-977 took beam from 4:50 PM to 6 PM; they had been off due to HP3DS work. NuMI resumed taking beam at 5 PM. Operations established store 7899 at 7:33 PM with an AIL of 328.2E30. A TeV person investigated a separator reversing switch problem that occurred during shot setup. MI personnel conducted slow spill studies.

Friday June 18

The midnight shift began with Operations monitoring store 7899 (148.34E30) and the stash (139.8E10), with Pbar stacking (25.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7899 at 4:03 AM, put the TeV into shot setup, and established store 7900 at 5:17 AM with an AIL of 248.1E30. Operations disabled beam to Booster, MI, and Recycler and prepared them for accesses.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/18/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory